what is the value of the expression 2x2+3xy-4y2 when x=2 and y=-4?
Found 2 solutions by RiverBaggett, mananth:
Answer by RiverBaggett(1)

2(x)2+3(x)(y)-4(y)2
2(2)2=3(2)(-4)-4(-4)=16

Answer by mananth(16946)

2x2+3xy-4y2 when x=2 and y=-4?
2*(2^2) +3(2)(-4)-4(-4)^2
=-80